An investment advisor invested $14,000 in two accounts. One investment earned 8% annual simple interest and the other investment earned 6.5% annual simple interest. The amount of interest earned for 1 year was $1024. How much was invested in each account?
Answer by stanbon(48550)

Let amount invested at 6.5% be "x"
The amount invested at 8% is "14000-x"
Interest on 6.5% money is 6.5%x= 0.065x
Interest on 8% money is 8%(14000-x)=0.08(14000)-0.08x=1120-0.08x
EQUATION:
Int. + Int. = $1024
0.065x+1120-0.08x=1024
-.015x= -96
x= 6400 dollars (amount invested at 6.5%
14000-6400=7600 (amount invested at 8%
Cheers,
Stan H.